# Fluid mechanics, Applied mathematics, Mathematical modelling

### v2

@@ Abstract
A central obstacle to understanding the route to turbulence in wall-bounded flows is that the flows are composed of complex, highly fluctuating, and strongly nonlinear states. In the case of pipe flow, models have deepened our understanding of turbulent onset by providing valuable theory to complement experiments and simulations. In planar cases, the large-scale flows associated with transitional turbulence are considerably more complex than for pipes, limiting our ability to develop models and provide theoretical analyses for these cases.
We address this challenge here by deriving from the Navier-Stokes equations a simplified model for transitional turbulence in a planar setting. The Reynolds-averaged and turbulent-kinetic-energy equations are projected onto a minimal set of wall-normal modes and justified model closures are used for the Reynolds stresses and turbulent dissipation and transport. The model reproduces phenomena found at the onset of turbulence in planar shear flows, such as turbulent-laminar patterns (turbulent bands) oriented obliquely to the streamwise direction and large-scale flows associated with both stationary patterns and growing turbulent spots. We demonstrate the model’s utility by showing that patterns arise with decreasing Reynolds number via a linear instability of uniform turbulence and by deriving a selection criterion for the pattern orientation at onset.
@@ 1 Introduction
The route to turbulence in many wall-bounded shear flows is mediated by a fascinating regime in which turbulence cannot be sustained throughout the system; rather, it occurs intermittently within laminar flow [1, 2, 3, 4, 5]. While the laminar state is stable to small perturbations, strongly nonlinear patches of turbulence may be sustained via interactions with neighbouring regions of quiescent laminar flow [6, 7, 8, 9]. Figure 1 illustrates the broad characteristics of wall-bounded transition in prototypical geometries: pipes and planar shear flows. In pipe flow, the turbulent patches take the form of axially localized turbulent structures known as puffs. Their dynamics and interactions along the pipe are crucial to the intermittent state mediating transition. In planar flows, transition is mediated by turbulent bands, also known as turbulent stripes, tilted obliquely to the streamwise (horizontal) direction. As is clearly evident, transition via turbulent bands is spatially richer than transition via puffs.
@@
A vast amount is known about both cases, but our theoretical understanding of transition in pipe flow is superior because we are able to reproduce and analyse the transition scenario with simple models. These are systems of partial differential equations that depend only on the axial coordinate and time, based on the interaction between large-scale (coarse-grained) turbulence and mean flow. Although the first models based on turbulent mean-flow interaction in pipes appeared more than a decade ago [10], there has been no adequate extension of such models to the planar case, and this has hindered our theoretical understanding of the transition to turbulence in wall-bounded flow. We address this here by deriving from the Navier-Stokes equations a simplified model for transitional turbulence in a planar setting.
@@
Figure 1: Sketch of the basic transition scenario in many wall-bounded shear flows. Flow states are illustrated for a pipe and shear flow between parallel planes, with the inset depicting the geometry rotated into perspective. Black signifies turbulent regions, where the fluid motion is highly fluctuating, while white signifies quiescent regions of smooth laminar motion. Three regimes are encountered as a function of the Reynolds number. At low $Re$, laminar motion is the only asymptotic state available to the system. At large $Re$, turbulence can be triggered, and its asymptotic state is spatially uniform, meaning that it adopts the full symmetry afforded by the system. Separating these two extremes is an intermittent regime where, once triggered, turbulence develops into a spatio-temporally intermittent form, composed of turbulent puffs or oblique turbulent bands. Two band states are illustrated: sparse bands (at lower $Re$) and dense bands (at higher $Re$). The sketches are not to scale: the length scale $L$ of turbulent structures, whether puffs or bands, is much larger than the wall separation $\ell$, which sets the scale of turbulent fluctuations. Coarse-grained modelling exploits that $\ell/L\ll 1$.
@@
We expand on the introductory paragraphs, beginning with further details on the intermittent route to turbulence, specifically the boundaries of the intermittent regime. As the flow rate, or Reynolds number ($Re$), decreases, puffs become sparse, and turbulent bands take on a fragmented form until a critical point is reached below which turbulence is no longer sustained. Universal scaling laws associated with directed percolation (DP) have been established for such critical points in some flows [11, 12, 13, 14]. Although critical phenomena have received considerable attention, this is not the focus of our work. As $Re$ increases, laminar regions disappear, and puffs and bands give way to a uniform turbulent state, whose mean recovers the full symmetry of the flow geometry. Equivalently, a symmetry breaking of the uniform turbulent state occurs as $Re$ decreases. A substantial body of work has analyzed and characterized this transition in the planar case [15, 16, 17, 18, 19, 9, 20, 21], but the understanding remains incomplete. This is, in part, the focus of our work. Figure 1 is intended to illustrate only generic characteristics of wall-bounded transition. In practice, many variations of such shear flows exist, differing in geometry (e.g., ducts, concentric cylinders, and disks) and in the mechanisms driving the shear (e.g., pressure gradients, moving walls, and body forces). While analogues of puffs and bands are generic, the details of intermittent state and what takes place at the boundaries of the intermittent regime may vary with the particular flow configuration.
@@
In this work we focus on a modelling approach aimed at describing the large-scale (coarse-grained) dynamics of the flows. As illustrated in figure 1, turbulent structures vary over distances much longer than the constrained direction of shear – the scale at which turbulence is generated [15, 22, 23, 24, 25]. Turbulent fluctuations may be viewed as “microscopic” [26, 27] and averaged over. For pipe flow, models in two scalar fields – the amplitudes of large-scale turbulence and mean flow – have proven successful in representing the dynamics of transitional turbulence in a single spatial coordinate (the pipe axis) and time [10, 28, 27]. Phenomenological equations capture turbulent puffs as smooth, non-fluctuating structures and provide an underlying bifurcation sequence for the route to turbulence, which can be readily analysed using dynamical systems theory [27, 29]. The role of “microscopic” turbulent fluctuations can be examined separately by incorporating noise terms into the model [26, 27]. Other modelling approaches to pipe flow exist. Notable cases are models based on azimuthal zonal flow [30], which have been extended to include effects of the large-scale streamwise flow [31], and models of puffs as discrete objects whose dynamics are modelled from experiments and simulations [32].
@@
Turbulent bands are more complex than turbulent puffs, and pipe models do not readily extend to the planar case. As a result, our understanding of bands is less developed than that of puffs. For example, turbulent bands are tilted relative to the streamwise flow direction, with typical angles ranging from approximately $20^{\circ}$ to $45^{\circ}$, depending on $Re$ and details of the flow geometry. We lack a fundamental mechanism explaining why bands are tilted and what determines their angle. This is not to say that there is no theoretical understanding of the band angle. Numerical studies have imposed, and thereby investigated the range of allowed tilt angles and how this depends on Reynolds number [7, 21]. and Kashyap et al. provide convincing, although indirect, evidence of a linear instability of the uniform turbulent state to turbulent bands tilted at $23^{\circ}\pm 0.5^{\circ}$ [19]. The origin of the band angle and the role of linear instability as a mechanism for symmetry breaking are precisely the types of issues that can be effectively addressed through modelling.
@@
The difficulty of extending pipe modelling arises primarily from the large-scale mean flow. In a pipe, the mean flow is effectively unidirectional and can be adequately represented by a single scalar field [27], whereas in the planar case, the mean flow must be described by a vector field (see figure 2b discussed below). Fluid advection plays a crucial role in these flows [7, 25, 9], and the vector nature of the mean flow must be properly accounted for if a model is to accurately capture the fluid mechanics of turbulent bands. There are alternative modelling approaches to planar flows distinct from that pursued in our work. One broad approach is to reduce the wall-normal resolution, either by lowering the resolution in an otherwise standard simulation [33] or by projecting the Navier–Stokes equations onto a few modal basis functions [34, 12, 35, 36, 37]. Such approaches can capture bands and the vector character of mean flow, but require the simulation of small-scale turbulence fluctuations. Another interesting modelling strategy introduces spatial coupling to the ordinary differential equations describing the self-sustaining process of wall-bounded turbulence [18, 38]. While these models provide insights into turbulent bands, they lack a vector representation of the large-scale flow and the associated mechanisms.
@@ 2.1 Geometry and governing equations
We derive our model from the most mathematically tractable planar flow that exhibits turbulent-laminar bands: a shear layer between parallel, stress-free boundaries, driven by a sinusoidal body force rather than by wall motion or a pressure gradient. Studies of this configuration in the inviscid case date back to at least Tollmien [39]. In the transition literature, the system is called Waleffe flow (Wf), after Waleffe who used it in his work on the self-sustaining process [40]. It reproduces the same transitional phenomena as plane Couette flow (pCf) and other planar wall-bounded shear flows [34, 4, 12], while the sinusoidal driving force and stress-free boundary conditions significantly simplify both analytical and computational treatments.
@@
We use coordinates $(x,y,z)$ aligned with the streamwise, wall-normal (vertical), and spanwise directions, respectively. See figure  2a. The components of the velocity $\tilde{\bf v}$ are denoted by $(\tilde{u},\tilde{v},\tilde{w})$. We non-dimensionalize using a length scale $h$ equal to the half gap between the stress-free walls, and a velocity scale $U$, equal to the maximum laminar velocity. Thus, time is non-dimensionalized by the advective timescale $h/U$. In non-dimensional coordinates, the walls are at $y=\pm 1$. We use $\beta\equiv\pi/2$ for the lowest vertical wavenumber allowed by the geometry.
@@
where $\tilde{p}$ is the pressure, $Re\equiv Uh/\nu$ is the Reynolds number, and $\nu$ is the kinematic viscosity. In addition to pressure and viscosity, there are drag and body-force terms on the right-hand side of momentum equation (1). The drag term, $-\alpha\tilde{\bf v}_{H}$, only acts on (and against) the horizontal velocity components $\tilde{\bf v}_{H}\equiv\tilde{u}{\bf e}_{x}+\tilde{w}{\bf e}_{z}$. The parameter $\alpha$ is the drag coefficient and is taken to be small. Such a drag term, often called Rayleigh or Ekman friction, is used frequently in hydrodynamic modelling contexts to approximate the effect of friction due to a no-slip boundary that has been omitted from the model [41, 4, 12, 42, 43]. It was introduced by Chantry, et al.[12] in Wf and is only a significant source of dissipation relative to viscous effects for $y$-independent modes and large horizontal scales.
@@
This force drives a simple laminar flow of the form $\tilde{\bf v}_{lam}=\sin(\beta y){\bf e}_{x}$. Although the laminar solution has an inflection point, it is nevertheless linearly stable due to wall confinement [40].
@@
To establish the model closures and calibrate the parameter values introduced below, we performed direct numerical simulations (DNS) of the governing equations using the open-source pseudo-spectral code Dedalus [44]. We simulated the equations in a domain of dimensions $L_{x}\times 2\times L_{z}$, imposing periodic boundary conditions in the horizontal directions. We used a Crank-Nicolson, forward Euler time-stepping scheme with a time step $\Delta t=0.003$. A Fourier-spectral method with 3/2 dealiasing was applied in the horizontal directions, with a resolution of eight grid points per spatial unit. In the vertical direction, we used a SinCos basis with the same spatial resolution. We initialized the simulations with random conditions at higher $Re$, and we obtained the banded solution shown in figure 2b by incrementally decreasing Re from the resulting uniform state.
In addition to the DNS in streamwise–spanwise domains shown in figure 2b, we performed simulations of turbulent bands using the tilted domain approach introduced by Barkley & Tuckerman [23]. In this setup, the streamwise direction is shortened, the spanwise direction is extended, and the domain is then tilted at a fixed angle ($24^{\circ}$ in our case) so that the shortened direction aligns with the along-band direction (figure 2b). Averaging such simulations in the along-band and time directions gives profiles depending on the across-band direction only. These are presented in figures throughout this section and are used for final model calibration.
@@ 2.2 Reynolds Average
[…]
Following standard procedures [47], we arrive at the Reynolds-averaged equations of motion for the large-scale flow,

### v3

@@ Abstract
A central obstacle to understanding the route to turbulence in wall-bounded flows is that turbulence initially appears only intermittently within an otherwise laminar background. In the case of pipe flow, models have deepened our understanding of turbulent onset by providing valuable theory to complement experiments and simulations. In planar configurations, the large-scale flows associated with intermittent transitional turbulence are considerably more complex than for pipes, limiting our ability to develop models and provide theoretical analyses for these cases.
We address this challenge here by deriving from the Navier-Stokes equations a simplified model for transitional turbulence in a planar setting. The Reynolds-averaged and turbulent-kinetic-energy equations are projected onto a minimal set of wall-normal modes and justified model closures are used for the Reynolds stresses and turbulent dissipation and transport. The model reproduces phenomena found at the onset of turbulence in planar shear flows, such as turbulent-laminar patterns (turbulent bands) oriented obliquely to the streamwise direction and large-scale flows associated with both stationary patterns and growing turbulent spots. We demonstrate the model’s utility by showing that patterns arise with decreasing Reynolds number via a linear instability of uniform turbulence and by deriving a bound on the pattern angle at onset: $0<|\theta_{c}|<45^{\circ}$.
@@ 1 Introduction
The route to turbulence in many wall-bounded shear flows is mediated by a fascinating regime in which turbulence cannot be sustained throughout the system; rather, it occurs intermittently within laminar flow [manneville2015transition, manneville2016review, manneville2017laminar, tuckerman2020patterns, avila2023transition]. While the laminar state is stable to small perturbations, strongly nonlinear patches of turbulence may be sustained via interactions with neighbouring regions of quiescent laminar flow [wygnanski1973transition, barkley2007mean, song2017speed, gome2023patterns1]. Figure 1 illustrates the broad characteristics of wall-bounded transition in prototypical geometries: pipes and planar shear flows. In pipe flow, the turbulent patches take the form of axially localized turbulent structures known as puffs. Their dynamics and interactions along the pipe are crucial to the intermittent state mediating transition. In planar flows, transition is mediated by turbulent bands, also known as turbulent stripes, tilted obliquely to the streamwise (horizontal) direction. As is clearly evident, transition via turbulent bands is spatially richer than transition via puffs.
@@
A vast amount is known about both cases, but our theoretical understanding of transition in pipe flow is superior because we are able to reproduce and analyse the transition scenario with simple models. These are systems of partial differential equations that depend only on the axial coordinate and time, based on the interaction between large-scale (coarse-grained) turbulence and mean flow. Although the first models based on turbulent mean-flow interaction in pipes appeared more than a decade ago [barkley2011simplifying], there has been no adequate extension of such models to the planar case, and this has hindered our theoretical understanding of the transition to turbulence in wall-bounded flow. We address this here by deriving from the Navier-Stokes equations a simplified model for transitional turbulence in a planar setting.
@@
Figure 1: Sketch of the basic transition scenario in many wall-bounded shear flows. Flow states are illustrated for a pipe and shear flow between parallel planes, with the inset depicting the geometry rotated into perspective. Black signifies turbulent regions, where the fluid motion is highly fluctuating, while white signifies quiescent regions of smooth laminar motion. Three regimes are encountered as a function of the Reynolds number. At low ReRe, laminar motion is the only asymptotic state available to the system. At large ReRe, turbulence can be triggered, and its asymptotic state is spatially uniform, meaning that it adopts the full symmetry afforded by the system. Separating these two extremes is an intermittent regime where, once triggered, turbulence develops into a spatio-temporally intermittent form, composed of turbulent puffs or oblique turbulent bands. Two band states are illustrated: sparse bands (at lower ReRe) and dense bands (at higher ReRe). The sketches are not to scale: the length scale $L$ of turbulent structures, whether puffs or bands, is much larger than the wall separation $\ell$, which sets the scale of turbulent fluctuations. Coarse-grained modelling exploits that $\ell/L\ll 1$.
@@
We expand on the introductory paragraphs, beginning with further details on the intermittent route to turbulence, specifically the boundaries of the intermittent regime. As the flow rate, or Reynolds number (ReRe), decreases, puffs become sparse, and turbulent bands take on a fragmented form until a critical point is reached below which turbulence is no longer sustained. Universal scaling laws associated with directed percolation (DP) have been established for such critical points in some flows [lemoult2016directed, chantry_universal, klotz2022phase, hof2023directed]. Although critical phenomena have received considerable attention, this is not the focus of our work. As ReRe increases, laminar regions disappear, and puffs and bands give way to a uniform turbulent state, whose mean recovers the full symmetry of the flow geometry. Equivalently, a symmetry breaking of the uniform turbulent state occurs as ReRe decreases. A substantial body of work has analyzed and characterized this transition in the planar case [prigent2002large, tuckerman2011patterns, philip2011temporal, manneville2012turing, kashyap2022linear, gome2023patterns1, gome2023patterns2, reetz2019exact], but the understanding remains incomplete. This is, in part, the focus of our work. Figure 1 is intended to illustrate only generic characteristics of wall-bounded transition. Many variations of such shear flows exist, differing in geometry (e.g., ducts, concentric cylinders, and disks) and in the mechanisms driving the shear (e.g., pressure gradients, moving walls, and body forces). While analogues of puffs and bands are generic, the details of intermittent state and what takes place at the boundaries of the intermittent regime may vary with the particular flow configuration.
@@
In this work we focus on a modelling approach aimed at describing the large-scale (coarse-grained) dynamics of the flows. As illustrated in figure 1, turbulent structures vary over distances much longer than the constrained direction of shear – the scale at which turbulence is generated [prigent2002large, prigent2003long, barkley2005computational, duguet2010formation, duguet2013oblique]. Turbulent fluctuations may be viewed as “microscopic” [pomeau2015transition, barkley2016theoretical] and averaged over, an approach pioneered by Reynolds [Reynolds1895]. For pipe flow, models in two scalar fields – the amplitudes of large-scale turbulence and mean flow – have proven successful in representing the dynamics of transitional turbulence in a single spatial coordinate (the pipe axis) and time [barkley2011simplifying, barkley2015rise, barkley2016theoretical]. Phenomenological equations capture turbulent puffs as smooth, non-fluctuating structures and provide an underlying bifurcation sequence for the route to turbulence, which can be readily analysed using dynamical systems theory [barkley2016theoretical, frishman2022dynamical]. The role of “microscopic” turbulent fluctuations can be examined separately by incorporating noise terms into the model [pomeau2015transition, barkley2016theoretical]. Other modelling approaches to pipe flow exist. Notable cases are models based on azimuthal zonal flow [shih2016ecological], which have been extended to include effects of the large-scale streamwise flow [wang2022stochastic], and models of puffs as discrete objects whose dynamics are modelled from experiments and simulations [lemoult2024directed].
@@
Turbulent bands are more complex than turbulent puffs, and pipe models do not readily extend to the planar case. As a result, our understanding of bands is less developed than that of puffs. For example, turbulent bands are tilted relative to the streamwise flow direction, with typical angles ranging from approximately $20^{\circ}$ to $45^{\circ}$, depending on ReRe and details of the flow geometry. We lack a fundamental mechanism explaining why bands are tilted and what determines their angle. This is not to say that there is no theoretical understanding of the band angle. Numerical studies have imposed, and thereby investigated the range of allowed tilt angles and how this depends on Reynolds number [barkley2007mean, reetz2019exact]. Furthermore, using a series of numerical simulations, Kashyap et al. provide evidence of a linear instability of the uniform turbulent state to turbulent bands tilted at $23^{\circ}\pm 0.5^{\circ}$ [kashyap2022linear]. The origin of the band angle and the role of linear instability as a mechanism for symmetry breaking are precisely the types of issues that can be effectively addressed through modelling.
@@
The difficulty of extending pipe modelling arises primarily from the large-scale mean flow. In a pipe, the mean flow is effectively unidirectional and can be adequately represented by a single scalar field [barkley2016theoretical], whereas in the planar case, the mean flow must be described by a vector field (see figure 2b discussed below). Fluid advection plays a crucial role in these flows [barkley2007mean, duguet2013oblique, gome2023patterns1], and the vector nature of the mean flow must be properly accounted for if a model is to accurately capture the fluid mechanics of turbulent bands. There are alternative modelling approaches to planar flows distinct from that pursued in our work. One broad approach is to reduce the wall-normal resolution, either by lowering the resolution in an otherwise standard simulation [manneville2011modelling] or by projecting the Navier–Stokes equations onto a few modal basis functions [chantry2016turbulent, chantry_universal, seshasayanan2015laminar, lagha2007largescaleflow, lagha2007modeling]. Such approaches can capture bands and the vector character of mean flow, but require the simulation of small-scale turbulence fluctuations. Another interesting modelling strategy introduces spatial coupling to the ordinary differential equations describing the self-sustaining process of wall-bounded turbulence [manneville2012turing, kashyap2025laminar]. While these models provide insights into turbulent bands, they lack a vector representation of the large-scale flow and the associated mechanisms.
@@ 2.1 Geometry and governing equations
We derive our model from the most mathematically tractable planar flow that exhibits turbulent-laminar bands: a shear layer between parallel, stress-free boundaries, driven by a sinusoidal body force rather than by wall motion or a pressure gradient. Studies of this configuration in the inviscid case date back to at least Tollmien [tollmien1935]. In the transition literature, the system is called Waleffe flow (Wf), after Waleffe who used it in his work on the self-sustaining process [waleffe1997self]. It reproduces the same transitional phenomena as plane Couette flow (pCf) and other planar wall-bounded shear flows [chantry2016turbulent, tuckerman2020patterns, chantry_universal], while the sinusoidal driving force and stress-free boundary conditions significantly simplify both analytical and computational treatments.
@@
[…]
@@
where $\tilde{p}$ is the pressure, $Re\equiv Uh/\nu$ is the Reynolds number, and $\nu$ is the kinematic viscosity. In addition to pressure and viscosity, there are drag and body-force terms on the right-hand side of momentum equation (1). The drag term, $\alpha\tilde{\bf v}_{H}$, only acts on (and against) the horizontal velocity components $\tilde{\bf v}_{H}\equiv\tilde{u}{\bf e}_{x}+\tilde{w}{\bf e}_{z}$. The parameter $\alpha$ is the drag coefficient and is taken to be small. Such a drag term, often called Rayleigh or Ekman friction, is used frequently in hydrodynamic modelling contexts to approximate the effect of friction due to a no-slip boundary that has been omitted from the model [suri2014drag, tuckerman2020patterns, chantry_universal, marcus1998model, pedlosky2013geophysical]. It was introduced by Chantry et al. [chantry_universal] in Wf and is only a significant source of dissipation relative to viscous effects for $y$-independent modes and large horizontal scales. Since the goal of the model is to capture flow features on large horizontal scales, Ekman drag plays an important role. As part of our forthcoming model closure, the drag coefficient will ultimately take different values for the $y$-independent and $y$-dependent modes (§2.6).
@@
This force drives a simple laminar flow of the form $\tilde{\bf v}_{lam}=\sin(\beta y){\bf e}_{x}$. Although the laminar solution has an inflection point, it is nevertheless linearly stable due to wall confinement [waleffe1997self].
@@
To establish the model closures and calibrate the parameter values introduced below, we performed fully-resolved direct numerical simulations (DNS) of the governing equations using the open-source pseudo-spectral code Dedalus [burns2020dedalus]. We simulated the equations in a domain of dimensions $L_{x}\times 2\times L_{z}$, imposing periodic boundary conditions in the horizontal directions. For time-stepping, we used a Crank-Nicolson method for the linear terms and forward Euler for the nonlinear terms (‘CNAB1’ solver in Dedalus [burns2020dedalus]), with a time step $\Delta t=0.003$. A Fourier-spectral method with 3/2 dealiasing was applied in the horizontal directions, with a resolution of eight grid points per spatial unit. In the vertical direction, we used a SinCos basis with the same spatial resolution. We initialized the simulations with random conditions at higher ReRe, and we obtained the banded solution shown in figure 2b by incrementally decreasing Re from the resulting uniform state.
In addition to the DNS in streamwise–spanwise domains shown in figure 2b, we performed simulations of turbulent bands using the tilted domain approach introduced by Barkley & Tuckerman [barkley2005computational]. In this setup, the streamwise direction is shortened, the spanwise direction is extended, and the domain is then tilted at a fixed angle ($24^{\circ}$ in our case) so that the shortened direction aligns with the along-band direction (figure 2b). Averaging such simulations in the along-band and time directions gives profiles depending on the across-band direction only. These are presented in figures throughout this section and are used for final model calibration.
@@ 2.2 Reynolds Average
We apply Reynolds averaging to the momentum equation by letting $\tilde{\bf v}={\bf u}+{\bf u^{\prime}}$, where ${\bf u}\equiv\langle\tilde{\bf v}\rangle$ is the mean velocity field and $\bf u^{\prime}$} is the fluctuating velocity field satisfying $\langle{\bf u^{\prime}}\rangle=0$[Reynolds1895]. The components of $\bf u$} are denoted by $(u,v,w)$, and similarly ${\bf u}^{\prime}=(u^{\prime},v^{\prime},w^{\prime})$. We will refer to $\bf u$} as the large-scale flow. We define $q\equiv\langle{\bf u^{\prime}}\cdot{\bf u^{\prime}}\rangle/2$ and refer to this as the turbulent kinetic energy (TKE), although it should be understood that the fluctuating field $\bf u^{\prime}$} may be very small in some regions, reflecting small fluctuations about the laminar state. We are agnostic regarding the specific averaging operation $\langle\cdot\rangle$ used in the Reynolds averaging to defined the large-scale fields. In practice, for the flows of interest, ensemble averaging, spatio-temporal averaging, or filtering may be applied, e.g. [kashyap2022linear, gome2023patterns1, barkley2007mean, Ohnishi_2011, tuckerman2020patterns, duguet2013oblique, wang2020quadrupolar]. The averaging used below for model calibration is an along-band and time average [barkley2007mean, Ohnishi_2011, tuckerman2020patterns].
Following standard procedures [pope2000turbulent], we arrive at the Reynolds-averaged equations of motion for the large-scale flow,